The area of a rectangle is 21 square meters. Find the length and width of the rectangle if its length is 4 meters greater than its width. Use an equation and the formula for the Area of a rectangle = (width)(length).
Found 2 solutions by benazir.sj@gmail.com, solver91311:
Answer by benazir.sj@gmail.com(109)

Let us assume the width as x so and length as 4+x
given area =21 sqm
area =length *width
21= (4+x)* x
21=4x+x^2
we can change this eqaution into standard form
x^2+4x-21=0
now factor this eqaution
x^2+4x-21=0
x^2+7x-3x-21=0
x(x+7)-3(x+7)
(x-3)(x+7)=0
x=3,-7
since wehave to calculte the area we have to consider only the positive value of x
x=3

checking
x=3
21=(x+4)*x
21 =3+4*3
21=7*3
21=21
x=-7
21 =(-7+4)*3
21 =-3*3
21 notequal to -9

Answer by solver91311(24713)



Let represent the width. Then represents the length. Since area is length times width and the area is 21 square meters, in the first place it should be casual to the most obvious observer that the length and width have to be 7 and 3 meters, but you can always say:





And just solve the quadratic for

John

My calculator said it, I believe it, that settles it